Unreal Designs System Status

Page Loaded: 12:18:16 PM 29th June 2017 GMT    
Will refresh every 30 seconds. Time now: 12:18:16

  • Hosting System Domains

    Click to Show/Hide

  • Customer? Access the client control panel here.
  • Server Speed Graphs

    Click to Show/Hide

System Stability

Incident Log


11th April 2017   #0012

Power Restored 02:55AM

Grid power has been restored and 04.ws12r2.west-eu.deployment.unreal-designs.co.uk is now back online. Internal services only were affected.

Power Outage affecting 04.ws12r2.west-eu.deployment.unreal-designs.co.uk 02:30AM

A power outage has occurred, taking 04.ws12r2.west-eu.deployment.unreal-designs.co.uk offline. This appears to be a national grid outage. No client services are affected.

Declared as Minor Incident


1st April 2017   #0011

Hosting Upgrade Completed 11:00AM

All hosting systems have now been upgraded. No downtime was observed and all services are stable.

Hosting System Upgrade 10:30AM

We are upgrading our hosting system today. Systems should only be affected for short periods. We will keep you updated as we go.

Declared as Minor Incident


27th January 2017   #0010

Attack Mitigated, Services Fully Resumed 05:30PM

All services have now been fully restored and all attacks mitigated. We apologise to anyone who had their service effected.

Monitoring Services, Attack Mitigated 04:00PM

Our major host has now mitigated the attacks. We are monitoring our services to ensure they remain stable.

Attack Ongoing, Services Stable 12:00PM

The attack is continuing and is now being mitigated. Services remain accessible and stable through our mirrors.

Declared as Minor Incident

Major DDoS Attack on Hosting Provider 05:00AM

Our systems have detected a large scale DDoS attack on our main hosting provider. Mirrored services sholuld remian live.

Declared as Major Incident


1st December 2016   #0009

Stream Stable and Live 09:58AM

The HLS stream is stable and live, viewable via the event site or direct via the stream api.

HLS Stream Restarted 09:38AM

The encoding system has been restarted, hopefully should reduce issues.

HLS Stream Unstable 09:32AM

Issues have been noticed with regards to the quality and consistency of the HLS Stream.

Live Stream Services Put Live 09:00AM

Our HLS stream and the live event site have been put live worldwide.

Declared as Minor Incident


18th July 2016   #0008

CDN Mirrored Services Restored 02:29PM

CDN Services fix is working, all mirrored systems live and functioning.

CDN Mirrored Services Fix in Place 02:13PM

A fixed has been issued to the effected services. We are now monitoring performance.

Mirrored CDN Services Encountering Issues 02:06PM

Issues have been reported on the initial CDN mirrored services.

Mirrored Systems Put Live 02:00PM

Our mirrored systems for Client Systems, Database Systems & CDN Services are now live.

Declared as Minor Incident


11th - 12th June 2016   #0007

Migration Completed 12th 11:25PM

All Systems migrated to new servers, all online and functioning.

SSL Connectivity Issues Fixed 12th 06:15PM

All SSL Connections now working. All sites live, awaiting IP transfer.

Declared as Minor Incident

Database Migration Completed 12th 01:45PM

All Databases migrated, awaiting final checks and IP transfer.

Free Month Hosting Rewarded 12th 12:00PM

Auto: Major systems have been down for 1 day, all clients rewarded with 1 month free hosting.

Error with Database Importing 12th 11:30AM

Unknown issue occured with Database importing, team investigating now.

Database Migration in Progress 11th 04:00PM

Database migration started, eta 09:00PM.

Site Migration Completed 11th 02:30PM

All sites being migrated now, database migration starting soon.

Site Migration in Progress 11th 12:30PM

All sites being migrated now, database migration starting soon.

Client Sites Offline for Maintenance 11th 12:00PM

Production servers taken offline for planned migration.

Declared as Major Incident


21st - 25th May 2016   #0006

Upgrades Completed 25th 07:43AM

All issues mitigated following upgrades, all systems functioning and online.

Data Transfer in Progress 24th 10:23PM

Hardware all installed, transferring all data.

Free Month Hosting Rewarded 24th 09:57AM

Auto: Major systems have been down for 1 day, all clients rewarded with 1 month free hosting.

Hardware Upgrade in Progress 23rd 09:57AM

Harddrive failure occured, replacements arrived, installing now.

Declared as Major Incident

Issues being Mitigated 22nd 10:14AM

Issues reported and located, awaiting fix decision from senior tech team.

Monitoring Upgrades 22nd 06:30AM

Instability reported, monitoring sites and upgrades.

Declared as Minor Incident

Upgrades Completed 21st 11:32AM

Upgrades all completed and all sites back online

Client Sites Offline for Maintenance 21st 09:45AM

Production servers & client sites taken offline for planned upgrades.

Declared as Major Incident


24th March 2016   #0005

Maintenance Completed 02:12PM

Following maintenance, the production servers are back online again.

Upgrades in Progress 02:09PM

New site styling and fixes being added. Soon to be completed.

Production Servers Offline for Maintenance 02:00PM

Production servers taken offline for planned maintenance.

Declared as Major Incident


23rd January 2016   #0004

Maintainence Completed 08:23AM

Following maintenance, the backup system is back online again.

Maintenance in Progress 08:16AM

Hourly database backups initialised, OnApp restarting.

Maintenance in Progress 08:05AM

Maintenance in progress on the backup system. Updates will be added as needed.

Backup Systems Offline for Maintenance 08:00AM

Backup system taken offline for planned maintenance.

Declared as Minor Incident


22nd September 2015   #0003

Fatal Error Fixed 02:36PM

Server updated and rebooted, production servers are back online.

Error being Investigated 01:02PM

Team investigating situation. Hope to have servers online soon.

Fatal Error on Production Servers 10:13AM

Production servers are down, error unknown, team responding asap.

Declared as Major Incident


19th September 2015   #0002

Maintenance Completed 03:17PM

Following maintenance, the backup system is back online again.

Maintenance in Progress 03:00PM

Maintenance in progress on the backup system. Updates will be added as needed.

Backup System Offline for Maintenance 02:58PM

Backup system taken offline for planned maintenance.

Declared as Minor Incident


28th August 2015   #0001

SPS Fixed and Online 01:51PM

Parts delivered and installed. Secondary power system back online.

Fix in Progress 09:24AM

New parts required and ordered. Service will continue with this system offline.

Issue being Investigated 09:02AM

Team on site, system failure being investigated. Updates will be added as needed.

Secondary Power System Failed 08:00AM

Secondary power supply has failed, reported to maintenance team.

Declared as Minor Incident